A buffer is made having a weak acid concentration of 0.50 M and a conjugate base concentration of 0.67 M. If the weak acid has a Ka of 4.5 x 10-6, what is the pH of the buffer? Answer with 3 sig figs.
Answer #1

Answer

pH = 5.48

Explanation

pKa = -logKa

pKa of the weak acid = -log(4.5×10-6) = 5.35

Henderson - equation is

pH = pKa + log([A-]/[HA])

pH = 5.35 + log(0.67M/0.50M)

pH = 5.35 + 0.13

pH = 5.48